OpenStudy (anonymous):

Dave travels 7 miles north then 24 miles west. How far is he from his starting point?

OpenStudy (anonymous):

@Falco276 and @thomaster

OpenStudy (falco276):

Does this include a coordinate graph?

OpenStudy (anonymous):

No

OpenStudy (espex):



OpenStudy (anonymous):

That's it Pythagorean Theron

OpenStudy (espex):

Then you know that \[\sqrt{7^{2}+24^{2}}=x\]

thomaster (thomaster):

\(7^2+24^2=x^2\) so the square root of this